OpenStudy (anonymous):

1. If a and b are any two real numbers, explain how you calculate the difference a – b. Be sure to include the following in your explanation: a. A description of the relationship between addition and subtraction b. How you can use a number line to help with the calculation c. How you can determine the sign of the result @samsungfanboy @sur

OpenStudy (samsungfanboy):

You got the mighty freckles on your post ;)

OpenStudy (anonymous):

Haha @SamsungFanBoy

OpenStudy (freckles):

How about we do a few examples? And you make your own summary from the examples? example to compute 5-3 you get 2  (this above picture is a number line showing how to do 5-3) to compute 3-5 notice 3>5 and you are subtracting a bigger number from a smaller number so to make this easier you can write this as -(5-3) and use the difference above we found to compute 5-3 and then get the result for -(5-3) as -2 now say we have an example of taking a negative from a positive 5-(-3) well this is the same as saying 5+3 since the opposite of a negative is positive 5-(-3)=5+3=8 say you wanted to take away a negative from a negative -5-(-3) well again-(-3) is still +3 so you have -5+3 which is the same as -(5-3) which is -2